Cathy and Ernie had four consecutive even integers. They found that the product of the second and fourth was 16 less than the product of -3 and the sum of the first and third. What were their integers?
Answer by Alan3354(69443)

Cathy and Ernie had four consecutive even integers. They found that the product of the second and fourth was 16 less than the product of -3 and the sum of the first and third. What were their integers?
------------
Use n-2, n, n+2 and n+4
---
n*(n+4) + 16 = -3*((n-2)+(n+2))
n^2 + 4n + 16 = -6n
n^2 + 10n + 16 = 0
(n + 2)*(n + 8) = 0
n = -2 ----> -4, -2, 0, +2 (If zero is considered an even integer)
==================
n = -8 ----> -10, -8, -6, -4